OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at L (-4, 0), M (4, 6), N (7, 2), O (-1, -4). Round your answer to the nearest hundredth when necessary.

OpenStudy (wolf1728):

LO² = MN² = 4² + 3² = 25 LO = MN = 5 LM² = ON² = 8² + 6² = 100 LM + ON = 10 So, perimeter = 2*5 + 2*10 = 30